A sample of 60 Grade 9 students’ ages was obtained to estimate the mean age of all Grade 9 students. =15.3 years and the population variance is 16.


1.) What is the point estimate for ?

2.) Find the 95% confidence interval for .

3.) Find the 99% confidence interval for .

4.) What conclusions can you make based on each estimate?


1
Expert's answer
2021-05-31T16:24:09-0400

1)

The sample mean is a point estimate of the population mean.


2)



"-1.96<\frac{\overline{x}-\mu}{\sigma/\sqrt{n}}<1.96"

For sample mean:

"\mu-1.96\sigma/\sqrt{n}<\overline{x}<\mu+1.96\sigma/\sqrt{n}"

"15.3-1.96\cdot4/\sqrt{60}<\overline{x}<15.3+1.96\cdot4/\sqrt{60}"

"14.29<\overline{x}<16.31"


3)

For the 99% confidence interval:

"Z_{0.95}=\pm2.58"

"-2.58<\frac{\overline{x}-\mu}{\sigma/\sqrt{n}}<2.58"

"15.3-2.58\cdot4/\sqrt{60}<\overline{x}<15.3+2.58\cdot4/\sqrt{60}"

"13.97<\overline{x}<16.63"


4)

Both intervals lie within one "\sigma" about the mean of population, and the 99% confidence interval is larger than the 95% confidence interval for the sample mean.